OpenStudy (anonymous):

Find the quadratic equation in x whose roots are 2a + 1/b and 2b + 1/a, given that ab = 3.5, a + b = -2.

OpenStudy (anonymous):

\[Ax^2+Bx+C=0\] sum of the roots is \(-\frac{B}{A}=-2\) product of the roots is \(\frac{C}{A}=3.5\) perhaps that will help is it \(2a+\frac{1}{b}\) or \(\frac{2a+1}{b}\)?

OpenStudy (anonymous):

the former

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

how do I continue?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

but what i wrote is wrong anyway a and b are not the roots, so that is a mistake maybe another tactic is needed

OpenStudy (anonymous):

ok i have an idea lets multiply the two roots and see what we get

OpenStudy (anonymous):

No you are right, the roots are 2alpha + 1/beta, and 2beta + 1/alpha, I couldn't find the symbols for alpha and beta so i used a and b

OpenStudy (anonymous):

\[(2a+\frac{1}{b})(2b+\frac{1}{a})=4ab+2a+2b+\frac{1}{ab}\] if my algebra is right

OpenStudy (anonymous):

no my algebra is bad tonight, let me try it again

OpenStudy (anonymous):

\[(2\alpha+\frac{1}{\beta})(2\beta+\frac{1}{\alpha})=4\alpha \beta+4+\frac{1}{\alpha\beta}\] i think that is better

OpenStudy (anonymous):

Ok got it. Thanks!

OpenStudy (anonymous):

we know all those numbers because \(\alpha \beta=\frac{7}{2}\) if my arithmetic is correct, that number is \(\frac{128}{7}\) but maybe my arithmetic is as bad as my algebra

OpenStudy (anonymous):

you good from there? you still have to add them

OpenStudy (anonymous):

i get the sum of the roots is \(-\frac{32}{7}\) but you should check it

OpenStudy (anonymous):

i like this question, it is cute